OpenStudy (anonymous):

Solve the equation. Check your answer. 3x – 7 = 5x + 19

OpenStudy (anonymous):

-2x = 26 x = =13 check. 3(-13) - 7 = 5(-13)+19 -39 - 7 = -65 + 19 -46 = - 46

OpenStudy (anonymous):

x = -13 ***